OpenStudy (anonymous):

medal will be given help 1 question r^2 - 2r -15

OpenStudy (anonymous):

I think this what you trying to do right? r2 + -2r + -15 Reorder the terms: -15 + -2r + r2 Factor a trinomial. (-3 + -1r)(5 + -1r) Final result: (-3 + -1r)(5 + -1r)

OpenStudy (anonymous):

Oh my bad, my bad. Don't listen to what I just did this is it! Step 1 : Simplify r2-2r - 15 Trying to factor by splitting the middle term 1.1 Factoring r2-2r-15 The first term is, r2 its coefficient is 1 The middle term is, -2r its coefficient is -2 The last term, "the constant", is -15 Step-1 : Multiply the coefficient of the first term by the constant 1 • -15 = -15 Step-2 : Find two factors of -15 whose sum equals the coefficient of the middle term, which is -2 -15 + 1 = -14 -5 + 3 = -2 Tiger found it!! Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, (-5) and (3) r2 - 5r + 3r - 15 Step-4 : Add up the first 2 terms, pulling out like factors : r • (r-5) Add up the last 2 terms, pulling out common factors : 3 • (r-5) Now add up the four terms of step 3 : (r+3) • (r-5) Which is the desired factorization Final result : (r + 3) • (r - 5)

OpenStudy (anonymous):

r^2-2r-15 1*15=-15 5-3=2 5*-3=-15

OpenStudy (anonymous):

OMGGGGGGG THANK YOU SO MUCHHHHHHHHHHHHH CAN YOU HELP ME WITH ONE MORE

OpenStudy (anonymous):

thank you for the answer also @surjithayer

OpenStudy (anonymous):

1. 6x^2 -5x-6 2. v=8y^3 + 10y^2 +3y last ones promise!!!

OpenStudy (anonymous):

@surjithayer That was the right answer?

OpenStudy (anonymous):

yessss that was the right but they put it in R + 3r -5 but you had it correct!!!

OpenStudy (anonymous):

1. 6*-6=-36 9*-4=-36 9-4=5 write -5=-(9-4) 2. v=y(8y^2+10y+3) 8*3=24 6*4=10 6*4=24

OpenStudy (anonymous):

1. 6x^2 -5x-6 Step 1 : Simplify 6x2-5x - 6 Trying to factor by splitting the middle term 1.1 Factoring 6x2-5x-6 The first term is, 6x2 its coefficient is 6 The middle term is, -5x its coefficient is -5 The last term, "the constant", is -6 Step-1 : Multiply the coefficient of the first term by the constant 6 • -6 = -36 Step-2 : Find two factors of -36 whose sum equals the coefficient of the middle term, which is -5 -36 + 1 = -35 -18 + 2 = -16 -12 + 3 = -9 -9 + 4 = -5 Tiger found it!! Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, (-9) and (4) 6x2 - 9x + 4x - 6 Step-4 : Add up the first 2 terms, pulling out like factors : 3x • (2x-3) Add up the last 2 terms, pulling out common factors : 2 • (2x-3) Now add up the four terms of step 3 : (3x+2) • (2x-3) Which is the desired factorization Final result : (2x - 3) • (3x + 2)

OpenStudy (anonymous):

that asnwer was correct and perfect dudeeeeeeeee.

OpenStudy (anonymous):

nevermind on the other one lollll

OpenStudy (anonymous):

@bettylee I'm sorry me or him? lol

OpenStudy (anonymous):

LOLLLLLL you ( :

OpenStudy (anonymous):

Lol No prob, glad I could help! ✌ and ❤